Minute of KAGRA commissioning meeting
- Date: 2018/10/2 11AM~12PM
- Attendance: Miyakawa, Yamamoto, Yokozawa, Kozakai, Kokeyama
- Zoom: Ushiba, Suzuki, Enomoto
X arm alignment
- Beam is returned back to central area from ETMX!
- Waiting for ITMX fixed.
- Oplev moved somehow. It is not clear why.
Trouble of ITMX
Trouble of PR2 and PR3
- Seems to be stable for PMC, IMC.
- 5W will be out in the PSL.
- New driver seemed to be not working. Miyo-kun tried it in other network, the it worked.
- Trouble might be network traffic in kAGRA control.
- EPICS gateway has been implemented. Network traffic should be much smaller now (~1/10).
- Still not perfect. Needs more tuning.
- Installation of periscope.
- Next meeting: 10/9.